OpenStudy (anonymous):

a parachutist exits a plane at a height of 4250ft directly over his target the path of his descent is a straight line that makes an angle of 85 degrees with the ground. by what distance does the parachutist miss his target?

OpenStudy (campbell_st):

start by drawing a picture  hope this helps.

OpenStudy (anonymous):

so use tangent

OpenStudy (campbell_st):

yep... thats the ratio so \[\tan(85) = \frac{4250}{d}\] solve for d

OpenStudy (anonymous):

so i multiply 85*4250=361,250 or is that wrong

OpenStudy (campbell_st):

nope when the unknown (d) is in the denominator just swap it with tan(85) so \[d = \frac{4350}{\tan(85)}\] and think carefully about you answer... does it make sense... you have the person missing by around 70 miles... so use the equation above to find the distance he misses by

OpenStudy (anonymous):

so 51.18ft

OpenStudy (anonymous):

so i dived 85/4350 or do i divide 4350/85

OpenStudy (campbell_st):

use the solution I'vw written above.